5. Stability of marriage. There are four men, α, β, γ, δ, and four women, A, B,C,D, The first number of each second number is the ranking of men by the women. For example, a-male ranks women C on third place, while C ranks α on the first place. A marriage profile is a pairing of men to women. A marriage profile is stable if there is no pair of a man and a woman who are not married to each other, and the woman prefers this man to her husband, and this man prefers this woman to his wife. Find a stable marriage profile.1234 3321 2341 2213 3412 1132 αβγδ
Answer #1

Preferences of men:

\alpha = A,B,C,D
\beta = A,B,C,D
\gamma = B,C,A,D
\delta = C,A,B,D

Preferences of women:

A = \gamma,\delta,\alpha,\beta
B = \delta,\alpha,\beta,\gamma
C = \alpha,\beta,\gamma,\delta
D = \delta,\gamma,\alpha,\beta

What does the algorithm look like? The algorithm is called "men-proposing algorithm". In each iteration, an unmarried man proposes to the first woman on his list that he hasn’t proposed to yet. A woman who receives a proposal that she prefers to her current assignment accepts it and rejects her current assignment.

  • Now, \alpha proposes to the first woman on his list: A. A has no offers and accepts.
  • \beta proposes to A, who rejects. He moves on to the second woman on his list: B. She has no offers and accepts.
  • Now, \gamma proposes to B. B prefers \beta over \gamma and rejects him. \gamma moves on to the second woman on his list: C. C has no offers and accepts.
  • Next, \delta asks out C, who refuses, because she prefers \gamma over \delta. \delta moves on to the second woman on his list: A. Now, A prefers \delta to \alpha, so she agrees and rejects \alpha.
  • \alpha moves on to B, who prefers him to \beta.
  • \beta asks out C who prefers him over \gamma, she accepts his offer and rejects \gamma.
  • \gamma asks out A, who prefers him over everyone else, she accepts \gamma and rejects \delta.
  • \delta asks out B, the third woman on his list, who prefers him over others, so she accepts him and rejects \alpha.
  • \alpha asks out C, who prefers him over everyone else, and she accepts him, rejects \beta.
  • \beta asks out the final woman on his list, D, who has no offers and accepts.

Hence, we get,
\alpha with C
\beta with D
\gamma with A
\delta with B
This association meets the definition of a stable marriage profile.

Hope this helped!